OpenStudy (anonymous):

Mitzi commutes 8 miles from her home to Valencia. What would an equivalent commute be in kilometers? Round off your decimal answer to 2 decimal places. Enter only significant zeros.

OpenStudy (anonymous):

@SalvadorV

OpenStudy (anonymous):

we know that 1 mile=1.609 kilometers, so you just have to make a rule of three. How would you do that?

OpenStudy (anonymous):

I have no idea.

OpenStudy (anonymous):

Round trip is 16 miles There are 1.61 miles per km. Multiply and prosper.

OpenStudy (anonymous):

what you do is, if 1 mile is 1.609 kilometers you need to know how much is 8 miles, so what you have to do is increase both by the same amount if 1 mi= 1.609 km then 2 mi= 3.218 km you can do that with any number, for example 5 1 mi = 1.609 km 1*5 mi =1.609*5km, so 5 miles = 8.045

OpenStudy (anonymous):

how would you do that with 8 miles?

OpenStudy (anonymous):

I got 12.872

OpenStudy (anonymous):

@SalvadorV

OpenStudy (anonymous):

awesome! that is correct

OpenStudy (anonymous):

now to roud it to the decimal

OpenStudy (anonymous):

how would you round that as I told you?

OpenStudy (anonymous):

12.87 ?

OpenStudy (anonymous):

16 x 1.61 = 25.76 km round -trip commute 8 x 1.61 = 12.88 one-way

OpenStudy (anonymous):

yup, 12.87 :D

OpenStudy (anonymous):

wait but he said its 12.88 ???

OpenStudy (anonymous):

so which one.

OpenStudy (anonymous):

@SalvadorV @douglaswinslowcooper

OpenStudy (anonymous):

he rounded it from the beginning, we rounded at the end, the result is a bit different, but I believe your teacher won't mind, it's up to you which one you prefer

OpenStudy (anonymous):

its not my teacher its an online quiz so now which one is a better choice.lol

OpenStudy (anonymous):

I'm not sure, is it multiple choice?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

Did not realize your situation one-way d = 8 x 1.609=12.872 rounds to 12.87 if start with 1.61 km/mi, then you get my 12.88

OpenStudy (anonymous):

I guess both can be right :D